OpenStudy (anonymous):

Find all x such that the distance between (x,-10) and (-20,-5) is the square root of 425.

OpenStudy (anonymous):

Distance formula: \[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] \[\sqrt{(x+20)^{2}+(-10+5)^{2}}=\sqrt{425}\] \[(x+20)^{2} + (-10+5)^2\] = 425 (x+20)^2 = 400 x+20 = +/-20 x = 0,-40